### v3

# Cobblestone: A Divide-and-Conquer Approach for Automating Formal Verification

Abstract.
Formal verification using proof assistants, such as Coq, is an effective way of improving software quality, but requires significant effort and expertise. Machine learning can automatically synthesize proofs, but such tools are able to prove only a fraction of desired software properties. We introduce Cobblestone, a divide-and-conquer approach for proof synthesis. Cobblestone uses a large language model (LLM) to generate potential proofs, uses those proofs to break the problem into simpler parts, automatically identifies which of those parts were successfully proven, and iterates on the remaining parts to build a correct proof that is guaranteed to be sound, despite the reliance on unsound LLMs. We evaluate Cobblestone on four benchmarks of open-source Coq projects, controlling for training data leakage. Fully automatically, Cobblestone outperforms state-of-the-art non-LLM tools, and proves many theorems that other LLM-based tools cannot, and on many benchmarks, outperforms them. Each Cobblestone run costs only \$1.25 and takes 14.7 minutes, on average. Cobblestone can also be used with external input, from a user or another tool, providing a proof structure or relevant lemmas. Evaluated with such an oracle, Cobblestone proves up to 58% of theorems. Overall, our research shows that tools can make use of partial progress and external input to more effectively automate formal verification.
copyright: none
1. Introduction
Bugs in software systems can be costly and dangerous. In 2022, poor software quality cost the US economy \$2.41 trillion (Krasner22, ), and bugs can bring down critical, global systems (Menn24, ). Formal verification using proof assistants, such as Coq (coq, ) or Lean (Moura21, ), is a promising method of improving software quality. Formal verification can mathematically prove the absence of entire classes of bugs, providing strong guarantees for the correctness of critical software systems. And formal verification is highly effective: A study (Yang11a, ) of C compilers found bugs in every tested compiler, including LLVM (Lattner04, ) and GCC (Stallman12, ), but not in the formally verified (in Coq) portions of CompCert (Leroy09, ).
But, formal verification requires specifying desired properties, writing mathematical proofs of the properties, and machine checking those proofs using the proof assistant. Writing these proofs requires significant expertise and manual effort. For example, the proofs verifying CompCert are 8 times longer than the functional code (Leroy06, ), and even small changes to the software can require heavy proof editing (Ringer21, ). While hundreds of large software systems have been verified (Ringer21, ), like the sel4 microkernel (Klein09, ; Murray13, ) and the CakeML compiler (kumar2014cakeml, ), and formal verification has seen industrial success, e.g., at Airbus France (Souyris14, ), Google, and Mozilla (Erbsen19, ; Jacobs20, ), most software today is not verified due to the high manual cost.
Recent work has aimed to reduce the cost of formal verification by using machine learning to synthesize proofs (Sanchez20, ; First20oopsla, ; First22icse, ; Sanchez-Stern23toplas, ; Paliwal20, ; Yang19, ; First23fse, ). However, these approaches, even when combined with an SMT-solver-based hammer (Czajka18, ) can only prove one third of the desired properties on a large benchmark of open-source Coq projects (First22icse, ). The advent of large language models (LLMs) has improved performance in this space (lu2024proofautomationlargelanguage, ; thompson2025rango, ; thakur2024an, ). In this paper, we aim to occupy the happy medium between two common approaches to proof synthesis.
Some work uses LLMs to generate whole proofs at once, without using the intermediate proof state produced by partial proofs (First23fse, ; zhang2023getting, ). For instance, PALM (lu2024proofautomationlargelanguage, ) samples an LLM for a whole proof attempt and then uses repair mechanisms. The success of whole-proof-based approaches relies on the powerful reasoning abilities of proprietary LLMs, such as OpenAI and Anthropic models, as they must construct proofs that follow viable high-level plans. However, these approaches fail to provide the LLM with access to the intermediate state, which contains important context to aid in reasoning.
Other approaches perform tactic-by-tactic search, building a proof one step at a time and allowing the LLM to condition the next step generation based on access to the intermediate state (Yang19, ; First20oopsla, ; Sanchez20, ; Yang23, ; Blaauwbroek24ICML, ; Jiang2022Thor, ). However, navigating the complex search space of possible proofs is challenging since the local state steers the search without consideration for a high-level proof plan.
These approaches represent two extremes on the spectrum of proof synthesis granularity. Our work is a middle ground between these two extremes, incorporating the advantages of whole-proof generation with leveraging intermediate states where this generation fails. To achieve this middle ground, we make a fundamental insight: we can leverage the tree structure of whole-proof attempts to localize valid and invalid subproofs. We can keep valid subproofs and recurse on points of failure where we regain access to the state. This insight motivated the design of our tool Cobblestone, a novel LLM approach to proof synthesis.
We present Cobblestone, which uses a divide and conquer algorithm to combine an LLM’s high-level planning ability with the advantages of local exploration. Cobblestone starts by tackling the entire goal, automatically identifying which parts of its proof are working and which are not, progressively reducing the granularity of its attack from the whole proof to individual proof steps by recursing, if necessary.
Figure 1. Given a software property to prove, Cobblestone first attempts to Refer to caption use CoqHammer to prove that property. If that fails, it Refer to caption generates a set of whole proofs, Refer to caption localizes errors in those proofs, Refer to caption extracts the working parts, and Refer to caption recurses on the remaining unproven subgoals to assemble correct proofs. “API access points” provide external information as discussed in Section 3.5
Cobblestone uses an LLM to generate whole proofs, analyzes the proofs using its novel fail-safe execution to decompose the proof into subgoals, and determines which subgoals’ subproofs work and which fail. It then attempts to repair the failing subproofs, and, if necessary, recurses on the remaining, unproven subgoals to build a whole, correct proof. Because Cobblestone divides the goal into smaller subgoals, it is more effective at using automated theorem proving tools such as hammers; while a hammer may fail to prove the overall goal, it is more likely to succeed on smaller, decomposed goals. Unlike prior work POETRY (wang2024proving, ), Cobblestone does not require fine-tuning an LLM to identify proof substructure and locations to recurse, making it easily extensible to use different LLMs (requiring only minor changes to API calls).
We implement Cobblestone for the Coq proof assistant, and evaluate it on GPT-4 and two other LLMs. We evaluate on subsets of four benchmarks: CoqGym (Yang19, ), a set of open-source Coq projects from GitHub used for evaluating prior proof synthesis tools (Sanchez20, ; First20oopsla, ; First22icse, ; Sanchez-Stern23toplas, ; Yang19, ); and coq-wigderson (10.1145/3618305.3623600, ), coq-bb5 (bb5github, ), and PnVRocqLib (pnvgithub, ), three recent projects deliberately chosen to be outside of the time range of GPT-4’s training cutoff to control for the impact of pretraining data leakage on LLMs.
We compare Cobblestone to prior non-LLM-based proof synthesis tools, LLM-based tools, and LLM baselines of our own creation. Prior RNN-based proof synthesis tool, Proverbot9001 (Sanchez20, ) proves 17% of the CoqGym subset and 10% of the coq-wigderson subset, while SMT-solver-based CoqHammer (Czajka18, ) proves 30% and 27%, respectively. We implement a baseline LLM-based approach that uses chain-of-thought reasoning and show that it proves 25% and 19%, respectively. Meanwhile, Cobblestone, automatically proves 48% of the CoqGym subset and 38% of the coq-wigderson subset. When prior LLM-based tool PALM (lu2024proofautomationlargelanguage, ) is run multiple times to match the token usage of our approach, Cobblestone outperforms it on the coq-wigderson, coq-bb5, and PnVRocqLib subsets, and is highly complementary with it on the CoqGym subset. Therefore, Cobblestone complements prior work by proving different theorems they do not, so that together, they prove significantly more.
While completely automated, Cobblestone can also use external input, such as proof structure plans or relevant lemmas from other tools or a proof engineer, further improving proof-synthesis success. Combining automated Cobblestone with versions that have access to a proof structure and relevant lemmas can prove 58% of the CoqGym subset and 55% of the coq-wigderson subset.
The main contributions of our work are:
  • Cobblestone, a novel LLM-based, divide-and-conquer proof-synthesis approach that uses partial successes of failing proofs to produce whole, correct proofs. This serves as a middle-ground strategy between whole-proof and tactic-by-tactic search.
  • A method for executing a proof to localize proof errors and separate working and failing parts of a proof.
  • A Cobblestone evaluation on four Coq benchmarks and a comparison to state-of-the-art proof synthesis tools showing that Cobblestone consistently proves theorems that prior work cannot, often outright proving more than them.
  • A data-based exploration of how external input from proof engineers or other tools can further improve Cobblestone’s success as part of interactive, semi-automated proof-synthesis.
To ensure reproducibility of our results and enable others to build on our work, we will make all code, experimental scripts, and data publicly available (anonCobblestoneReplicationPackage, ).
2. The Cobblestone Approach
To formally verify a software system, in addition to writing the code for the system, an engineer needs to write a mathematical theorem formalizing the desired property, and a high-level proof (proof script) that the property holds. Cobblestone automatically writes the proof script. Given code and a theorem stating a property (e.g., in_adj_exists in Figure 2), Cobblestone tries to synthesize a proof using a divide and conquer approach, as shown in Figure 1. First, Cobblestone tries synthesizing a proof using CoqHammer (Section 2.3). If that fails, Cobblestone samples an LLM for candidate solutions (Section 2.4). If none of the candidates prove the theorem entirely, Cobblestone uses a novel fault localization method to isolate errors while noting which subgoals are successfully proven (Section 2.5). Next, Cobblestone keeps successful parts of candidate solutions and recurses to find proofs for the yet unproven subgoals (Section 2.6), ultimately combining them into a proof of the original theorem.
While Cobblestone is fully automated, it has extension points (labeled “API access point for external information” in Figure 1) that can be used to provide useful additional information. Section 3.5 will evaluate how this external information can increase Cobblestone’s proving power.
2.1. Illustrative Example
1Definition adj (g: graph) (i: node) : nodeset :=
2 match M.find i g with
3 | Some a => a
4 | None => S.empty end.
5
6Lemma in_adj_exists : forall g i j,
7 S.In i (adj g j) ->
8 exists v, M.find j g = Some v /\ S.In i v.
9Proof.
10 intros g i j H.
11 unfold adj in *.
12 destruct M.find eqn:E in *.
13 - hammer.
14 - rewrite SP.FM.empty_iff in H. contradiction.
15Qed.
Figure 2. A sample theorem of the in_adj_exists property and its human-written proof (lines 9–15).
We illustrate our approach using a real-world theorem in_adj
_exists
.111from graph.v in the coq-wigderson project Figure 2 lists a definition (lines 1–4), the in_adj_exists theorem statement (lines 6–8), and a human-written proof for the theorem (lines 9–15). This theorem states that for all graphs g and all nodes i, j, if i is adjacent to j in g, then g contains v, a set of nodes adjacent to j, and that i $\in$ v.
Coq proofs consist of high-level commands called tactics, like “try induction on n” or “try simplifying.” Tactics transform the proof state, which contains goals (statements that must be proven) and assumptions that can be used to prove the goals. The proof state starts with the initial theorem to be proven. When tactics transform the proof state to contain no more goals, the Qed command completes the proof.
Figure 3(a) shows the proof state before executing the “destruct M.find eqn:E in *” tactic on line 12 of Figure 2, which performs case analysis. The proof state consists of a single goal and several assumptions, including H, which states that i is a member of adj g j. [^2] The unfold tactic on line 11 inlines adj in the proof state. Executing the tactic results in the proof state in Figure 3(b). The new proof state decomposes the preceding single goal into two subgoals. Each subgoal replaces “M.find” in the preceding goal with one of its two return values — “Some v” or “None”. If both subgoals are proven, then the original theorem is true.
g: graph
i: S.elt
j: node
H: S.In i (
match M.find j g
with
| Some a => a
| None => S.empty
end)
(1/1)
exists v : nodeset,
M.find j g = Some v /\
S.In i v
g: graph
E: M.find j g = Some n
H: S.In i n
(1/2)
exists v : nodeset, Some n = Some v
/\ S.In i v
H: S.In i S.empty
(2/2)
exists v : nodeset, None = Some v
/\ S.In i v
Figure 3. The proof states (a) before and (b) after executing destruct on line 12 of Figure 2.
The rest of the proof script is split into two subproofs, one for each subgoal, each starting with a bullet (-). The first bullet on line 13 invokes CoqHammer using the hammer tactic. CoqHammer (Czajka18, ) attempts to automatically prove a goal using SMT solvers and proof reconstruction procedures. If hammer succeeds, the goal is proven. Otherwise, the tactic produces an error.
While hammer is able to find a proof for the first subgoal, it has limitations. For example, CoqHammer cannot perform induction, a key step in many Coq proofs. In this case, hammer cannot prove the second subgoal. Instead, after the bullet on line 14 of Figure 2, the proof uses the rewrite tactic, which takes advantage of an already proven lemma (SP.FM.empty_iff) to change the hypothesis H to enable a proof by contradiction. After executing the rest of this proof script, which successfully proves the theorem, the proof state will contain no more goals, indicating that the theorem is proven.
2.2. Cobblestone on the Illustrative Example
Figure 4. Proving in_adj_exists with Cobblestone. Refer to caption hammer fails, and an LLM-generated proof contains errors, resulting in recursive calls for each subgoal. Refer to caption The first subgoal is proven using a new LLM completion and hammer. Refer to caption The second subgoal is proven using hammer. Refer to caption The final correct proof.
We now demonstrate how Cobblestone proves in_adj_exists automatically. Figure 4 displays the steps Cobblestone takes to divide the proof into smaller pieces and prove each of them, resulting in a correct proof. First, Cobblestone attempts to prove in_adj_exists in one go, trying hammer, which fails, and then prompting an LLM to generate whole proofs, one of which is shown in [Uncaptioned image]. While the LLM’s proof starts off with reasonable structure, it contains errors (highlighted in red). Normally, Coq will not be able to identify all the errors; Section 2.5 describes Cobblestone’s novel fault localization mechanism. Though the proof in [Uncaptioned image] is not correct as-is, it does decompose the initial goal into two subgoals. It may still be a useful part of a correct proof if these subgoals can be proven. However, its correctness can only be determined retrospectively after proving these subgoals.
Instead of naïvely trying to prove in_adj_exists again, as prior work has tried (First23fse, ), Cobblestone recurses on each subgoal, attempting to prove it instead ([Uncaptioned image] and [Uncaptioned image]). In [Uncaptioned image], Cobblestone tries using a hammer, which fails, and then prompts an LLM for a proof of the subgoal “M.find j g = Some (adj g j)”. This results in a proof with an error, which Cobblestone temporarily fixes by adding the admit tactic. The proof with admit, a proof in progress, is passed as an argument to another recursive call. Here, Cobblestone changes the proof in progress to a working proof by replacing admit with the hammer tactic. In [Uncaptioned image], Cobblestone is able to prove “S.In i (adj g j)” using only hammer.
With successful subproofs for each subgoal in hand, Cobblestone combines parts of the proofs in [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image] to form a correct proof of the whole theorem, shown in [Uncaptioned image]. Notably, by using LLM completions and its divide-and-conquer approach, Cobblestone can divide and simplify the goals, often reaching subgoals simple enough for hammer to prove. This allows the final proof to integrate portions of two separate LLM completions (for different subgoals), and two invocations of hammer, into a single proof.
2.3. Applying CoqHammer
Cobblestone combines neural-based proof synthesis and SMT-based proof synthesis by invoking CoqHammer (Czajka18, ). In Figure 1, the hammer icon ([Uncaptioned image]) shows how Cobblestone uses CoqHammer in two ways. First, at the start of Cobblestone’s execution (and in subsequent recursive calls), Cobblestone invokes CoqHammer on the goal, which sometimes produces a working proof. Second, in recursive calls, Cobblestone may be passed a proof in progress, which consists of working tactics followed by an “admit”. For example, “unfold adj, M.find in *. admit.” in Figure 4 [Uncaptioned image] is a proof in progress. Cobblestone replaces admit with hammer, and checks if this proves the subgoal. Attempting to repair proofs in progress like this allows Cobblestone to use CoqHammer on a variety of goals, increasing the frequency of success.
2.4. Sampling LLM for Whole Proofs
Central to Cobblestone’s proof synthesis is using an instruction-tuned LLM (ouyang2022training, ) to sample whole proofs. Cobblestone uses a prompt that consists of two strings — a setup prompt with high-level directions about the task the LLM should perform, and a sampling prompt with details specific to the current theorem. In LLM terminology, the setup prompt is called a “system message” and the sampling prompt is called a “user message,” though Cobblestone generates these prompts automatically, without a user.
The setup prompt directs the LLM to produce a whole proof for the provided theorem. The sampling prompt contains 5 types of information: (1) the theorem statement to prove, (2) the current proof state, (3) the definitions for all identifiers mentioned in the theorem statement and proof state that are not in the standard library, (4) contextual information, and (5) optional reasoning, which is used as part of chain-of-thought prompting. The setup prompt is formatted as below, with each section preceded by a section header. Cobblestone constructs the sampling prompt fully automatically using the coq-serapy library (Sanchez20, ) to interact with Coq.
[CURRENT THEOREM]
{{theorem statement from Coq ITP}}
[PROOF CONTEXT]
{{current proof context from Coq ITP}}
[DEFINITIONS]
{{definitions from Coq ITP}}
[OTHER PROVEN THEOREMS]
{{proven theorems from Coq ITP and optional oracle}}
[REASONING]
{{optional chain-of-thought reasoning}}
The contextual information in the prompt can either be empty or contain the theorem statements that were successfully proven before the given theorem in the file. If the information in the prompt is longer than the LLM’s token limit, the contextual information is left-truncated, removing information from the file that is further from the theorem statement. Cobblestone samples the LLM with this prompt, which results in candidate whole proof scripts.
Chain-of-thought (Wei2022-qv, ), a popular prompting framework that prompts an LLM to provide natural language reasoning before generating its final response, has been shown to help improve LLM performance. Cobblestone implements chain-of-thought by first prompting the LLM with a modified setup prompt that instructs it to generate reasoning, and then prompting it a second time, with the generated reasoning added to the [REASONING] section of the sampling prompt. When Cobblestone samples an LLM without chain-of-thought prompting, this extra section is omitted.
Prior work has shown that the use of diverse data in models can increase the proving power of a proof-synthesis approach (First22icse, ). Accordingly, to increase variation in Cobblestone’s generated proofs, Cobblestone samples the LLM with 4 variations of the same prompt in 2 different dimensions — with and without contextual information and with and without chain-of-thought reasoning. Cobblestone benefits from this approach because the Coq theorem prover can serve as an oracle of proof correctness to pick a correct proof from among multiple samples. Our replication package (anonCobblestoneReplicationPackage, ) includes all prompts that were used as part of our evaluation.
2.5. Localizing Errors with Fail-Safe Mode
Figure 5. An execution of fail-safe mode on an input proof
Once Cobblestone samples proofs from the LLM, it executes each of them in Coq. If a proof proves the goal, Cobblestone returns it as-is. If none of the sampled proofs are successful, Cobblestone localizes the errors in each proof, identifying which parts are correct, and which parts are not. Executing a generated proof naïvely using the theorem prover stops on the first error, which then misses correct parts that occur after that error. Instead, we create fail-safe mode, a novel way of executing a proof in the theorem prover. This approach localizes errors by recursing over the proof’s structure, making small edits to allow the execution of proof sections even after an error, and annotating sections as “correct” and “in progress.” These annotations are implemented via a parallel data structure that keeps track of which subproofs are correct and which are in progress.
Fail-safe mode relies on the fact that many proofs (43% of proofs from the ChainOfThought evaluation from Section 3.2) contain subproofs, sections of the proof organized using bullets. A proof consists of a prefix followed by 0 or more subproofs, each of which consists of a bullet followed by a nested proof. The last tactic in the prefix decomposes a single goal into multiple subgoals, one for each subproof (recall Section 2.1). Each subproof’s bullet focuses Coq on a single subgoal in the proof state; the proof that follows proves that subgoal. Figure 6 formalizes this structure.
$\mathit{Proof}$ $\rightarrow$ $\mathit{Prefix}\ \mathit{Subproof}*$
$\mathit{Prefix}$ $\rightarrow$ $\mathit{Tactic}+$
$\mathit{Subproof}$ $\rightarrow$ $\mathit{Bullet}\ \mathit{Proof}$
[…]
Figure 6. The formal structure of a proof with bullets
Note that the definition of subproof is recursive. Each subproof can contain subproofs of its own (subsubproofs). In all figures, we indicate the level of nesting of the subproof by repeating the “-” character in bullets (i.e. “-” for a subproof, “--” for a subsubproof, etc.). A proof with no bullets can be expressed solely using a prefix and no subproofs. This means that all possible proofs can be expressed using this structure. (Section 3.6 will discuss fail-safe mode’s limitations in localizing errors in proofs consisting of only a prefix.)
Algorithm 1 Fail-safe mode
1:A prefix $p$, and 0 or more subproofs spssps
2:the modified proof runs without errors, and the most deeply nested subproofs with errors are annotated as “in progress”
3:procedure RunFailSafe($p$, spssps)
4:  for all tactics $t$ in $p$ do $\triangleright$ First, run $p$
5:   Run($t$)
6:   if $t$ results in an error then
7:     replace $t$ and everything after it with admit.
8:     annotate $p$ as “in progress”
9:      return
10:     annotate $p$ as “correct”
11:  for all subproofs $sp$ in spssps do $\triangleright$ then, run spssps
12:   RunFailSafe(spsp.prefix, spsp.subproofs)
Algorithm 1 describes fail-safe mode, which takes a proof script broken into prefixes and subproofs as input, modifies it to allow execution after errors, and annotates subproofs, subsubproofs, etc. as “correct” or “in progress.” First, it executes each tactic in the prefix. If a tactic results in an error, that tactic and everything after it is replaced with admit, and the prefix is annotated as “in progress.” Note that inserting admits into proofs in progress is temporary; Cobblestone will remove these in a recursive call.
If the prefix executes without an error, Cobblestone marks it as correct and fail-safe mode recurses on each of the subproofs, annotating them in the same way. Figure 5 shows the result of running fail-safe mode. Tactics that run successfully are colored green, tactics that produce errors are colored red, and admits are colored yellow as they are temporary. The boxes around prefixes and subproofs in the output are annotations, colored yellow if a section is in progress and green if it is correct. Note that Cobblestone marks the subsubproof starting with Tactic_E as in progress, but not the subproof starting with Tactic_D that contains it. Because annotations occur as part of a depth-first search, fail-safe mode localizes errors to the most deeply nested subproof possible.
In addition to the description in Algorithm 1, fail-safe mode also handles two complications that arise when executing LLM-generated proofs. To understand these complications, consider the example proof that an LLM could generate below and to the left.
Proof.
intros.
destruct a.
apply nonexistent.
- assumption.
- auto.
Qed.
Proof.
intros.
destruct a.
- admit.
- admit.
Qed.
The first complication is that the proof can encounter an error while executing a prefix tactic after it decomposes the original goal. Concretely, destruct a might successfully decompose the goal into subgoals, but then, apply nonexistent can fail. In this case, instead of just replacing the apply nonexistent with an admit, fail-safe mode searches for a shorter prefix that leads to multiple subgoals, which we call a decomposing prefix. For example, if the sequence intros. destruct a. leads to multiple subgoals, then fail-safe mode keeps it as the prefix and creates a bullet with an admit placeholder for each subgoal, as shown on the right.
The second complication is that a decomposing tactic can produce more or fewer subgoals than the number of bullets that follow. On the right, there are two bullets that follow destruct a. If destruct a produces fewer subgoals than two, fail-safe mode trims the number of bullets; if destruct a produces more than two, fail-safe mode adds the appropriate number of bullets with admit placeholders. This guarantees that the number of subproofs that appear syntactically in the proof matches the number of subgoals generated by the prefix.
2.6. Recursively Invoking Cobblestone
After executing fail-safe mode on each of the four sampled proofs, Cobblestone produces four annotated proofs like the one in Figure 5. This example contains two subproofs. The first one (starting with Tactic_D) has two subsubproofs. One is in progress, and the other succeeds. The second subproof is also in progress.
The overall theorem is not considered proven until Cobblestone generates and swaps in correct replacements for all the subproofs in progress for one of the four annotated proofs. To do this, Cobblestone loops over the annotated proofs. For each proof, Cobblestone makes one recursive call per subgoal in progress. It passes the subgoal and its corresponding subproof in progress (which contains admit) to the recursive call. Note that when fail-safe mode annotates a deeply nested subproof as in progress (e.g., the subsubproof that starts with Tactic_E), it only recurses on the corresponding subgoal, not on any of the subgoals containing it (e.g., the subgoal corresponding to Tactic_D). If none of the proofs are able to break the goal into subgoals, Cobblestone calls itself again, retrying with the original theorem as its goal. We limit this search procedure using two hyperparameters—maximum depth, which limits the level of recursion, and maximum invocations, which limits the total number of times Cobblestone can be invoked. These settings are discussed further in Section 3.1.
Through this recursion, Cobblestone uses divide-and-conquer, iteratively simplifying the goal and recursing on simpler subgoals, increasing the probability of success. This way, Cobblestone breaks down the problem of proving the original theorem into smaller chunks, and leverage progress made in the previous iterations to assist the subsequent ones. Cobblestone recurses in a greedy depth-first manner; preliminary experiments with backtracking and restarting from scratch do not improve performance.
3. Evaluation
We evaluate Cobblestone on four datasets of theorems from open-source Coq projects. We compare Cobblestone to five prior tools — CoqHammer, Proverbot9001, Tactician, PALM, and Rango— as well as two LLM baselines of our own creation.
Our evaluation answers five research questions:
  1. RQ1:
    How does Cobblestone compare to state-of-the-art proof synthesis methods?
  2. RQ2:
    How much does CoqHammer contribute to Cobblestone’s performance?
  3. RQ3:
    How much does Cobblestone’s search strategy contribute to its performance?
  4. RQ4:
    How does external information affect Cobblestone’s performance?
  5. RQ5:
    How do theorems Cobblestone proves and fails to prove differ?
3.1. Experimental Setup
Benchmarks. We construct our evaluation benchmarks from CoqGym’s test set of theorems (Yang19, ), and from three other open-source projects — coq-wigderson (10.1145/3618305.3623600, ), coq-bb5 (bb5github, ), and PnVRocqLib (pnvgithub, ).
CoqGym is a widely used benchmark for evaluating proof-synthesis tools (Yang19, ; First20oopsla, ; First22icse, ; Sanchez-Stern23toplas, ), comprised of 68,501 theorems and their associated human-written proofs across 124 projects. CoqGym’s test set consists of 26 projects with 12,161 theorems, including verdi-raft (distributed software correctness), goedel (Gödel’s 1st incompleteness theorem), and tree-automata (verified algorithms). Evaluating using this dataset allows for a more fair comparison to prior tools, and assessing tool efficacy across a diverse set of projects. However, CoqGym was released in 2019, and its projects existed on GitHub before then. Because the pre-training cutoff for GPT-4 is September 2021, GPT-4 has potentially seen CoqGym in its pre-training.
To address this concern, we also evaluate using coq-wigderson (10.1145/3618305.3623600, ), coq-bb5 (bb5github, ), and PnVRocqLib (pnvgithub, ). The first commit from each of these projects (March 2022, April 2024, and August 2024, respectively) is after the GPT-4 pre-training cutoff. coq-wigderson is a formal verification of Wigderson’s graph coloring algorithm. It consists of 174 theorems. PnVRocqLib contains 810 theorems about mathematics, including first order logic, Hilbert logic, and boolean algebra. coq-bb5 contains 1446 theorems about busy-beaver values.
Because of the cost of performing inference on LLMs, full-scale evaluations on tens of thousands of theorems are impractical. In fact, the cost of just the LLM usage for our evaluation, including ablation studies, exceeded US\$6K. Instead, we create four 100-proof subsets of the four benchmark projects, sampling theorems at random. We call these subsets CoqGym100, Wigderson100, PnVRocqLib100, and BB5100. To keep the cost of ablations low, we evaluate on PnVRocqLib100 and BB5100 only for RQ1 (Section 3.2). The total size of this dataset (400 theorems), is in line with other prior work which uses LLMs for proof synthesis (wanglego, ; wang2024proving, ; thakur2024an, ), and 300 of the 400 theorems are published after GPT-4’s training cutoff.
Comparisons to State-of-the-Art and Baselines. We compare Cobblestone to 5 prior proof-synthesis tools— (1) Proverbot9001 (Sanchez20, ), an RNN-based neural theorem prover; (2) CoqHammer (Czajka18, ), which uses SMT solvers and proof reconstruction procedures; (3) Tactician (Blaauwbroek20, ), which uses an online k-nearest-neighbor algorithm to select tactics during proof search; (4) PALM (lu2024proofautomationlargelanguage, ), which leverages retrieval augmentation (RAG), custom repair prompts, and a backtracking algorithm; and (5) Rango (thompson2025rango, ), which fine-tunes an LLM model and uses RAG that includes relevant lemmas and proof scripts. Proverbot9001 outperforms other ML-based proof synthesis tools that do not use LLMs, including ASTactic (Yang19, ), TacTok (First20oopsla, ), Diva (First22icse, ), and Passport (Sanchez-Stern23toplas, ). Proverbot9001 and CoqHammer prove 19.8% and 26.6% of theorems, respectively, on CoqGym, and prove 17% and 30% on CoqGym100. In a prior “informal comparison” of tools on a subset of CoqGym, Tactician outperformed Proverbot9001 (except on one project) and sometimes outperformed CoqHammer across projects (Blaauwbroek24ICML, ).
For CoqHammer, we use the Z3 (de2008z3, ), CVC4 (DBLP:conf/cav/BarrettCDHJKRT11, ), Vampire (10.1007/978-3-642-19835-9_7, ), and E (SCV:CADE-2019, ) SMT solvers, along with the default timeout settings (a 20 second prover timeout and a 5 second reconstruction timeout), following prior work (Sanchez20, ; Blaauwbroek24ICML, ). We evaluate Tactician’s ability to produce a proof fully automatically within a 10-minute timeout. Unlike Tactician, Proverbot9001 uses depth limits rather than a time limit, but usually finished its search within 10 minutes as well. We run PALM with gpt-4-0613 as its base model. One run of PALM on CoqGym100 used  3.7K tokens per theorem, 15 times less than the amount for Cobblestone (59K tokens per theorem). To correct for this, we run PALM 15 times for each dataset. Rango is run with the same setup noted in the paper—using a fine-tuned model with temperature 1.0 and a 10 minute timeout. Rango is run on an NVIDIA RTX 2080 GPU, along with a CPU with 16GB of RAM for proof checking.
We also compare against 2 new baselines, ChainOfThought, and LinearRegen. ChainOfThought prompts gpt-4-0613 20 times (the same as Cobblestone) in exactly the same way as Cobblestone does (using preceding lemmas and chain-of-thought prompting as described in Section 2.4) but only checks the proofs for correctness and does no further processing. LinearRegen also prompts gpt-4-0613 up to 20 times in exactly the same way as Cobblestone, but regenerates portions of a proof starting from its first failure point. When a proof has a failure, LinearRegen keeps the prefix of the proof that succeeds, and prompts the LLM to continue from the last successful proof state.
Metrics. In line with prior evaluations (First20oopsla, ; First22icse, ; First23fse, ; Sanchez-Stern23toplas, ), we use two metrics: success rate and added value. The success rate of a tool is the fraction of all theorems the tool is able to prove. The added value of tool X over tool Y is the number of new theorems X proves that Y does not, divided by the number of theorems Y proves. To measure the number of theorems a tool can prove in a fixed number of LLM samples, we define the proven@k metric. This metric is the sum of the pass@k (chen2021evaluating, ) values for each proof in a dataset, and represents the expected number of theorems proven with k samples.
Computing Resources. Throughout our evaluation, we use OpenAI’s GPT-4 model (gpt-4-0613), which has a context length of 8,192 tokens. In Section 3.2, we additionally use GPT-3.5-turbo (gpt-3.5-turbo-0125), and Claude 3 Opus (claude-3-opus-
20240229
). GPT-3.5-turbo has a context length of 16,385. Claude 3 Opus has a context length of 200,000 tokens, but to manage the cost of inference, we capped its token limit at 10,000. We chose these three models because their APIs support function calling, a key part of how our implementation samples whole proofs. We interface with GPT-3.5-turbo and GPT-4 through the OpenAI Python library. We use the official Anthropic library to interface with Claude 3 Opus. For each call to an LLM, we sample with temperature 1.0. All experiments are run on machines with Intel Xeon Gold 6230 CPUs and 125GB of memory. We have access to 10 cores from these CPUs, as our experiments were run in a virtualized environment.
We run Cobblestone with a maximum depth of 5, and invoke it up to 5 times. Because each Cobblestone invocation samples an LLM 4 times, Cobblestone uses up to 20 LLM samples in total.
3.2. RQ1: How Does Cobblestone Compare to State-of-the-Art Proof Synthesis Methods?
Figure LABEL:tab:rq1 shows the success rates for non-LLM tools (CoqHammer, Proverbot9001, and Tactician), our newly created baselines (ChainOfThought and LinearRegen), LLM-based tools (PALM and Rango), and Cobblestone. Note that Tactician and Rango cannot run on certain datasets (reported with N/A), as these datasets use unsupported older Coq versions.
Cobblestone proves 48% of theorems on CoqGym100 and 38% of theorems on Wigderson100. It outperforms each of the non-LLM baselines as well as the union of these baselines, All non-LLM, adding 54.5% additional value in CoqGym100, and 25.0%, 40.0% and 14.6% added value in Wigderson100, PnVRocqLib100, and BB5100 respectively. Across all datasets, Cobblestone proves 46 theorems All non-LLM cannot, All non-LLM proves 14 theorems that Cobblestone cannot, and both prove an overlap of 127 theorems.
On CoqGym100, running Cobblestone costs \$1.46 and takes in 15.32 minutes, with successful runs costing an average of \$0.23 and completing in 2.54 minutes. On Wigderson100, running Cobblestone costs \$1.03 and completes in 14.21 minutes, and successful runs cost on average \$0.15 and complete in 3.66 minutes.
Cobblestone proves more theorems than both ChainOfThought and LinearRegen in all datasets, adding a value of up to 100.0% over ChainOfThought, and up to 104.8% over LinearRegen. Cobblestone is also more token-efficient than LinearRegen. On CoqGym100 and Wigderson100, Cobblestone uses significantly fewer tokens (59.3K vs. 99.3K on CoqGym100 and 42.0K vs. 72.3K on Wigderson100). Similarly, on PnVRocqLib100, Cobblestone uses 51.8K, while LinearRegen uses 83.7K, and on BB5100, Cobblestone uses 47.8K while LinearRegen uses 80.6K.
Cobblestone also consistently proves theorems that the other LLM-based tools cannot in all datasets, outrightly proving more theorems than these tools in many cases. Cobblestone proves more theorems than a single run of PALM (PALM 1x in Figure LABEL:tab:rq1). It also proves more theorems than our PALM benchmark (which is run 15x) on Wigderson100, PnVRocqLib100, and BB5100, proving 8%, 9% and 2% more theorems (the difference in success rates) respectively. Furthermore, in all datasets, Cobblestone proves many theorems that PALM cannot. Cobblestone proves 46 theorems PALM cannot, PALM proves 32 theorems Cobblestone cannot, and both are able to prove an overlap of 126 theorems. On CoqGym100 and Wigderson100, Cobblestone uses a similar number of tokens to PALM, with an average of 59.3K vs. PALM’s 58.1K on CoqGym100 and 42.0K vs. PALM’s 41.8K on Wigderson100. On PnVRocqLib100, Cobblestone uses fewer tokens on average (51.8K vs. PALM’s 62.0K), while using more tokens on BB5100 (47.8K vs. PALM’s 40.7K). Cobblestone and PALM’s complementarity indicate that Cobblestone may benefit from using techniques such as PALM’s retrieval augmentation (RAG) and LLM-based proof repair.
Cobblestone proves the same number of theorems as Rango on BB5100, and significantly more on PnVRocqLib100. In both cases, Cobblestone and Rango also demonstrate significant complementarity. On BB5100, their union proves 55%, and on PnVRocqLib100, it proves 47%, more than either tool alone. Across all datasets, Cobblestone proves 29 theorems that Rango cannot, Rango proves 15 theorems that Cobblestone cannot, and they both prove an overlap of 55 theorems. This indicates that Cobblestone may also benefit from leveraging Rango’s RAG. Cobblestone and Rango’s token consumption is not straightforward to compare, as Rango uses a 1.3 billion parameter LLM, which is not directly comparable to gpt-4-0613, a much larger model.
Since Cobblestone uses an LLM as a black box, it works with any LLM that is fine-tuned for tool use. We modified Cobblestone and ChainOfThought to use GPT-3.5-turbo and Claude 3 Opus, as shown in Figure 7(b), and evaluated them on CoqGym100 and Wigderson100. For all three models, Cobblestone outperforms the combination of ChainOfThought and CoqHammer. However, using different LLMs affects performance relative to GPT-4. GPT-3.5-turbo underperforms, both in ChainOfThought and in Cobblestone. Claude 3 Opus outperforms GPT-4 when used in ChainOfThought, and has mixed results in Cobblestone, slightly overperforming on CoqGym100, and underperforming on Wigderson100.
CoqGym100
GPT-4 added GPT-3.5 added Claude added
value value value
CoqHammer 30% 66.7% 30% 040.0% 30% 66.0%
ChainOfThought 25% 92.0% 13% 200.0% 28% 78.6%
ChainOfThought 37% 35.1% 35% 025.7% 41% 24.4%
+CoqHammer
Cobblestone 48% 37% 49%
Wigderson100
GPT-4 added GPT-3.5 added Claude added
value value value
CoqHammer 27% 044.4% 27% 029.6% 27% 25.9%
ChainOfThought 19% 100.0% 08% 312.5% 22% 50.0%
ChainOfThought 31% 025.0% 28% 025.0% 31% 12.9%
+CoqHammer
Cobblestone 38% 33% 32%
(b) Figure 7b: Performance of Cobblestone and ChainOfThought implemented with different LLMs. “added value” is Cobblestone’s value added relative to each row.
RA1: Cobblestone outperforms non-LLM based proof generation methods and additional baselines. Cobblestone also performs comparably with LLM-based methods, complementing their results by proving theorems they cannot, and outperforming them in many cases. Cobblestone can be easily adapted to use LLMs other than GPT-4.
3.3. RQ2: How Much Does CoqHammer Contribute to Cobblestone’s Performance?
CoqGym100 Wigderson100
success added success added
rate value rate value
CoqHammer 30% 27%
ChainOfThought 25% 19%
ChainOfThought 37% 48.0% 31% 063.2%
$\cup$ CoqHammer
Cobblestone-NoHammer 25% 16%
Cobblestone-NoHammer 38% 52.0% 30% 87.5%
$\cup$ CoqHammer
Cobblestone 48% 96.0% 38% 137.5%
Figure 8. CoqHammer’s use contributes significantly to Cobblestone and ChainOfThought. “added value” is the value of adding CoqHammer.
Figure 8 shows that CoqHammer plays an important role. On CoqGym100, Cobblestone-NoHammer (an ablated version Cobblestone) only proves 25% of theorems, while Cobblestone proves 48%. On Wigderson100, it’s 16% vs. 38%. On the two benchmarks, CoqHammer helps Cobblestone prove 96.0% and 137.5% more theorems.
Even running CoqHammer once improves ChainOfThought’s success rate—from 22% to 36% for CoqGym100, and from 17% to 31% for Wigderson100. However, Cobblestone outperforms this simpler combination of the two methods. Recall that Cobblestone uses CoqHammer in each invocation. Another simple combination is to invoke CoqHammer once for each theorem, succeeding if it or Cobblestone-NoHammer succeeds. Cobblestone also outperforms this approach (“Cobblestone-NoHammer $\cup$ CoqHammer”). On CoqGym100, Cobblestone adds 96.0% value over Cobblestone-NoHammer, compared to the simpler approach’s 52%. On Wigderson100, Cobblestone adds 137.5%, vs. the simpler approach’s 87.5%.
RA2: CoqHammer and Cobblestone are significantly complementary. CoqHammer contributes to Cobblestone’s performance, but Cobblestone outperforms CoqHammer and other simple combinations of the two. Even when run without CoqHammer, Cobblestone proves theorems other methods cannot.
3.4. RQ3: How Much Does Cobblestone’s Search Strategy Contribute to Its Performance?
Wigderson100 CoqGym100
success added success added
rate value rate value
CoqHammer 27% 30%
ChainOfThought 19% 25%
Proverbot9001 10% 17%
Tactician 13% N/A
All Prior Together 34% 39%
TacticByTactic-NoHammer 08% 00.0% 11% 02.6%
Cobblestone-NoHammer 16% 00.0% 25% 07.7%
TacticByTactic 33% 11.8% 40% 20.5%
Cobblestone 38% 17.6% 48% 30.8%
Figure 9. Cobblestone outperforms a simpler search strategy, both with and without CoqHammer. “added value” is the value each tool adds over “All Prior Together”
Figure 10. Theorems proven vs. LLM samples. The y values for “ChainOfThought” and “CoqHammer+ChainOfThought” use proven@k, detailed in Section 3.1.
Cobblestone’s search strategy is influenced by three key design decisions which we will examine in this RQ—(1) its use of whole-proof completions (2) its use of a recursive divide-and-conquer approach, and (3) its use of a sampling temperature of 1.0.
Cobblestone’s use of whole-proof completions sets it apart from many other proof synthesis systems. By contrast, most neural theorem provers synthesize proofs using tactic-by-tactic search, predicting the next tactic and using a tree search (typically depth-first or beam search) to search the tactic space (Yang19, ; First20oopsla, ; First22icse, ; Sanchez-Stern23toplas, ; Sanchez20, ).
To measure the effectiveness of whole-proof completions, we implement another tool called TacticByTactic that, like Cobblestone, has use of the hammer, but uses GPT-4 to predict only the next tactic at each step of its search. While inefficient, TacticByTactic attempts the hammer tactic at each step of its search.
We run TacticByTactic, and TacticByTactic-NoHammer, a variant that does not use hammer, allowing up to 20 tactic predictions, with a maximum proof depth of 20, and 3 attempts to predict the next tactic at each step. As each prediction attempt results in one LLM sample, these tools use the same number of samples as Cobblestone.
Figure 9 shows that without using CoqHammer, TacticByTactic-NoHammer underperforms all prior tools on Wigderson100 and CoqGym100. With calls to CoqHammer, TacticByTactic performs much better, proving 33% and 40% of the theorems in Wigderson100 and CoqGym100, respectively, and outperforms Cobblestone-NoHammer. Still, Cobblestone outperforms TacticByTactic and adds more value both in CoqGym100 and Wigderson100.
The second distinguishing feature of Cobblestone’s search strategy is its use of a divide-and-conquer approach. As described in Section 2.6, when a proof does not work, Cobblestone invokes itself recursively to prove subgoals of the original theorem, rather than retrying on the theorem itself. Figure 10 shows how this recursive approach contributes to Cobblestone’s performance.
One important datapoint to observe is what happens with a single invocation (which only uses 4 samples). To observe this datapoint, one should look at Figure 10, at the 4 point on the X axis. Even with a single invocation, Cobblestone proves more theorems than the combination of CoqHammer and ChainOfThought can in 20 samples. This is thanks to Cobblestone’s error localization (Section 2.5) and careful application of CoqHammer (Section 2.3).
This datapoint is also interesting to observe because it is an ablation that runs fail-safe mode without recursively invoking Cobblestone. This ends up being similar to PALM’s (lu2024proofautomationlargelanguage, ) backtracking algorithm (which also does not make recursive calls).
As Cobblestone invokes itself recursively, using more LLM samples, the number of theorems it can prove increases even further. On CoqGym100, Cobblestone is able to prove 44 out of 48 (91.6%) theorems within 1 invocation, proving an additional 4 (9.4%) using recursive invocations. On Wigderson100, it is able to prove 31 out of 38 (81%) theorems within one invocation, proving 7 (19%) more over the course of its recursive calls. One possible explanation for this is test leakage in CoqGym100, which would make more samples of the original theorem “close enough” for Cobblestone’s to apply CoqHammer and error localization to produce a proof.
CoqGym100 Wigderson100
success added success added
rate value rate value
ChainOfThought-temp0 10% 9%
ChainOfThought 25% 150.0% 19% 111.1%
Cobblestone-temp0 48% 32%
Cobblestone 48% 000.0% 38% 025.0%
Figure 11. A temperature of 1 may improve performance, and does not harm performance. “added value” is value of changing from temperature 0 to 1.
The third distinguishing feature of Cobblestone’s search strategy is its use of temperature 1.0 when sampling whole proofs. Cobblestone benefits from sampling diverse proofs (recall Section 2.4). Setting the temperature to 1.0 contributes to this diversity. To examine the benefits of a higher temperature, we ran versions of Cobblestone and ChainOfThought with the temperature set to 0. Figure 11 shows the performance of these versions. In both Wigderson100 and CoqGym100, ChainOfThought-temp0 proves a subset of the theorems that ChainOfThought is able to prove. In Wigderson100, Cobblestone proves more theorems than Cobblestone-temp0, with a value added of 25%, and in CoqGym100, it proves the same number of theorems.
RA3: Cobblestone’s search strategy outperforms prior tools and an LLM-based tactic-by-tactic search. A single invocation of Cobblestone outperforms a combination of CoqHammer and ChainOfThought, and Cobblestone proves even more theorems when invoked recursively. A temperature of 1.0 can also improve performance.
3.5. RQ4: How Does External Information Affect Cobblestone’s Performance?
Wigderson100 CoqGym100
success added success added
rate value rate value
Cobblestone 38% 48%
Cobblestone-PerfPrems 43% 21.1% 50% 12.5%
Cobblestone-PerfDecomp 52% 42.1% 47% 14.6%
All 3 Cobblestone versions 55% 44.7% 58% 20.8%
Figure 12. External information in terms of relevant lemmas (Cobblestone-PerfPrems) and a breakdown of the proof into subgoals (Cobblestone-PerfDecomp) significantly improves Cobblestone’s proving power. “added value” is the value each tool adds over Cobblestone.
Cobblestone can use external information from another tool or a human. As a proxy, we created two oracles to provide such information (via the “API access points” in Figure 1. For each theorem, the perfect premises oracle (PerfPrems) knows the set of lemmas already proven in the project that the human-written proof for this theorem uses (e.g., with “rewrite” as described in Section 2.1). These lemmas are provided in addition to existing contextual information in the [OTHER PROVEN THEOREMS] section of the prompt in Section 2.4. Meanwhile, the perfect decomposition oracle (PerfDecomp) knows the set of subgoals the human-written proof proves. PerfDecomp provides a decomposing prefix (recall Section 2.5) from the human-written proof, helping fail-safe mode break the theorem into subgoals for recursive calls.
Using the PerfPrems oracle is equivalent to asking a human (or a tool) “What lemmas are relevant to proving this theorem?” and using the PerfDecomp oracle is equivalent to asking “How would you break down this theorem into smaller goals?”
We next evaluate Cobblestone with access to these oracles, calling the resulting variants Cobblestone-PerfPrems and Cobblestone-PerfDecomp. Cobblestone-PerfDecomp has access to both PerfPrems and PerfDecomp oracles. Figure 12 shows that Cobblestone-PerfPrems outperforms Cobblestone on both datasets, and Cobblestone-PerfDecomp sometimes performs even better. Interestingly, there is some complementarity to the variants, with each proving some theorems none of the others do. Together, the variants are able to prove 55% of CoqGym100 and 58% of Wigderson100. The additional information provided to Cobblestone by PerfDecomp does not always result in more theorems proven. Cobblestone-PerfPremises and Cobblestone are able to prove 6 theorems in CoqGym100 and 3 theorems in Wigderson100 that Cobblestone-PerfDecomp does not. This indicates that running Cobblestone several times, with different kinds of information in each run, may improve performance.
RA4: External information, such as useful already-proven lemmas or a decomposition of the theorem into subgoals can significantly increase Cobblestone’s proving power, suggesting a promising direction for future research.
3.6. RQ5: How do theorems Cobblestone proves and fails to prove differ?
To better understand where Cobblestone succeeds, we examine proofs that Cobblestone, Cobblestone-PerfPrems, and Cobblestone-PerfDecomp generate. To better understand when Cobblestone fails, we also examine the human-written proofs of some theorems where it fails. All proofs referenced in this section are available in our replication package(anonCobblestoneReplicationPackage, ).
Successes: Cobblestone’s successful proofs range from 1 to 24 tactics in length, with the shortest consisting of a single invocation to CoqHammer. Successful proofs with length greater than 1 have an average length of 8.3 tactics, with bullet depth ranging from 1 (no bullets) to 4 (3 levels of bullets). These proofs can consist of anywhere from 1 to 9 distinct CoqHammer invocations or LLM samples. Of the proofs that decomposed their goal using bullets, prefixes ranged from 1 to 4 tactics long, with an average of 1.6. Of these prefixes, 41% decomposed using split, 36% used induction, 18% used destruct, and 14% used apply.
The projects in our benchmarks broadly fall into two categories—mechanizations of mathematical proofs, and formalizations of software. CoqGym100 consists of 12 mathematical projects and 8 software projects, with 36 and 64 theorems in each category. Cobblestone proves 61.1% of CoqGym100’s mathematical theorems, and 40.6% of its software theorems. Additionally, coq-wigderson is a software project, and coq-bb5 and PnVRocqLib are mathematical projects. In total, our benchmarks contain 236 mathematical theorems and 164 software theorems, with Cobblestone proving 46.2% and 39.0% of these theorems respectively.
Together, Cobblestone-PerfPrems and Cobblestone-PerfDecomp generate 26 successful proofs for theorems unproven by Cobblestone (10 in CoqGym100 and 16 in Wigderson100). Interestingly, one of these proofs is of votesWithLog_update_elections_
data_timeout
,333from RefinementSpecLemmas.v in verdi-raft whose human-written proof relies on break_match and tuple_inversion—custom tactics in Coq’s Ltac language. Cobblestone-PerfPrems generates an alternate proof without custom tactics. In the proof of Mcardinal_Scardinal,444from graph.v in coq-wigderson an LLM sample asserts a proposition that helps prove the theorem.
Failures: Despite its improvements over the state-of-the-art, Cobblestone and its variants that use external information are unable to prove 45 theorems from Wigderson100 and 42 theorems from CoqGym100. We randomly sampled 14 such theorems (7 from each dataset) to examine manually.
Many of these theorems’ human-written proofs are significantly longer than those that our tools generated proofs for. Cobblestone’s average proof is 10 tactics long, whereas 5 of the 14 unproven theorems have a human-written proof 20 or more tactics long. We also found that some theorems make it difficult for Cobblestone to take advantage of internal progress. For example, the ground-truth proof of requestVoteReply_term_sanity_client_
request555from RequestVoteReplyTermSanityProof.v in verdi-raft
primarily consists of tactics like unfold and apply, each of which helps prove the theorem, with none breaking the goal into subgoals. It is difficult for Cobblestone to generate such proofs as it requires generating a working proof with a single LLM sample, preventing its divide and conquer algorithm from being able to recurse. Generalizing our approach to use partial progress that does not involve splitting a goal into subgoals may allow Cobblestone to prove more such theorems.
RA5: Cobblestone’s proofs are sometimes shorter than their human-written counterparts, and often successfully leverage its divide and conquer approach. Cobblestone’s failures indicate an opportunity to leverage new kinds of partial progress.
3.7. Threats to Validity
All LLM evaluations may suffer from test data leaking into the pretraining dataset, reducing generalizability to unseen data. We mitigate this risk by using coq-wigderson, coq-bb5, and PnVRocqLib, whose first GitHub commits are after GPT-4’s publicly stated pretraining cutoff date (recall Section 3.1). We evaluate on 400 theorems, in line with other published work (wanglego, ; wang2024proving, ; thakur2024an, ), to mitigate cost.
While our approach is general, our implementation is specific to Coq and uses GPT-4. We also evaluate with GPT-3.5-turbo and Claude 3 Opus to demonstrate generalizability.
Finally, our evaluation only began exploring how external information can aid Cobblestone’s proof synthesis and used oracles synthesized using human-written proofs. The intent of these oracles is to proxy how a human might interact with Cobblestone, but are too perfect to make any predictions on how an interactive, semi-automated approach may perform. More research is necessary to study the effects of incomplete and noisy data, as well as user studies to show the potential impact of real, human-provided information.
4. Related Work
Recent work in automating theorem proving in proof assistants has mostly explored three overarching approaches: hammers, machine learning techniques, and combinations of the two. Hammers, such as CoqHammer (Czajka18, ) and Sledgehammer (Paulson23, ) for Isabelle, call out to SMT solvers (schulz2013system, ; deMoura08, ) to construct low-level proofs, though they cannot use induction, and so are limited in what they can prove on their own. Our evaluation has shown that our approach can often prove theorems CoqHammer fails to prove on its own.
Machine learning techniques called neural theorem provers typically use a predictive model learned from a corpus of existing proofs to predict the next steps, and use these predictions to guide a search through the space of potential proofs (li2024survey, ). They have been built using RNNs (Huang18, ; Sanchez20, ), LSTMs (Yang19, ; First20oopsla, ; First22icse, ), GNNs (bansal2019learning, ; Blaauwbroek24ICML, ), and transformer-based LLMs (Jiang21, ; Yang23, ). LLM-based methods prompt the pretrained model zero-shot or with few-shot examples (jiang2023draft, ; zhang2023getting, ), fine-tune the model on proof data (First23fse, ; huang2024mustard, ), or use it as an agent (thakur2023languageagent, ; wanglego, ).
Hammers and machine learning techniques are complementary (Jiang2022Thor, ; First22icse, ). Thor (Jiang2022Thor, ) fine-tunes an LLM to learn when to apply Sledgehammer to a subgoal or when to predict a tactic. By contrast, Cobblestone samples whole proofs rather than tactics and makes hammer calls without prediction. Focusing on formalizing math, DraftSketchProve (jiang2023draft, ) uses an LLM to translate informal proofs into a formal proof sketch with holes, filling in the holes with calls to Sledgehammer, with others extending this framework (zheng2023lyra, ; wanglego, ; zhao2023decomposing, ). LEGO (wanglego, ) decomposes theorems into helper lemmas using an LLM, providing no guarantee that proving them will prove the original theorem; by contrast, Cobblestone uses the theorem prover to decompose goals with such a guarantee.
LeanDojo (Yang23, ) and Magnushammer (Mikula24, ) train models to select relevant premises (lemmas/definitions) at each proof step. In addition to selecting relevant premises, Rango (thompson2025rango, ) also selects similar proof scripts. Cobblestone uses the preceding lemmas in the file, but could benefit from a premise selection model or a proof-script selection model, in the same way it benefits from external information in the form of an oracle’s perfect premises.
Proof engineers often repair previously working proofs (Ringer20, ). Symbolic tools can automate some repair (Ringer21, ; masci2022proof, ). Baldur uses error messages to fine-tune an LLM to repair Isabelle proofs, in the context of proof synthesis (First23fse, ). Unlike Baldur, which only attempts one repair, Cobblestone is an iterative, divide-and-conquer approach. PALM (lu2024proofautomationlargelanguage, ) also samples from an LLM and uses repair mechanisms. However, as detailed in Section 3.4, our approach is recursive in nature.
POETRY fine-tunes an LLM and estimates proof points on which to recurse (wang2024proving, ). By contrast, Cobblestone requires no tuning training data and uses fail-safe mode to precisely identify subgoals that recursion. Adapting Cobblestone to use more powerful models requires only minor changes to its API calls, whereas adapting POETRY is only possible with access to the model’s weights, and would require significant resources for fine-tuning.
Formally specifying system properties is complementary to proving them. Prior work has tackled formalizing natural language specifications (Endres24, ; Goffi16, ; Motwani19icse, ; Zhai20, ). Once specified, some properties, like fairness (Galhotra17fse, ), can be verified probabilistically (Thomas19science, ; Hoag23icse-demo, ; Metevier19neurips, ; Giguere22iclr, ; Albarghouthi17, ).
Prompt engineering is an important aspect of effective LLM use (openai2023gpt4, ; chowdhery2023palm, ; touvron2023llama, ). Prompting frameworks, such as chain-of-thought (wei2022chain, ) and others (yao2023tree, ; chen2024boosting, ; besta2023graph, ), can elicit reasoning in LLMs (wei2022chain, ; wang2022self, ; zhou2022least, ; fu2022complexity, ; chu2023survey, ). LLMs can be used for quantitative reasoning tasks (Lewkowycz2022Minerva, ; azerbayev2023llemma, ) and programming tasks (rozière2023code, ; chen2021evaluating, ). These advances are complementary to our efforts to generate proofs automatically.
5. Contributions
Cobblestone is a novel divide-and-conquer method for synthesizing formal verification proofs using LLMs. Cobblestone significantly outperforms prior tools and LLM-based baselines, proves more complex theorems, and complements other LLM-based tools. Cobblestone demonstrates a promising potential for collaborating with humans or other tools to synthesize even more proofs.
6. Acknowledgements
This work is supported by the National Science Foundation (NSF) under grant no. CCF-2210243, NSF award no. 2107397, and by Defense Advanced Research Projects Agencies (DARPA) under Contract No. HR0011-24-2-0307. The hammer icon used in Figure 1 was provided by the Noun Project.